[ Five years ago Max was one third as old as his father. In 25 years he will be three fifths as old as his father. Find their ages now. ]
Answer by TimothyLamb(4379)

m - 5 = (1/3)(f - 5)
m = (1/3)(f - 5) + 5
m = (1/3)f - (5/3) + (15/3)
m = (1/3)f + (10/3)
m + 25 = (3/5)(f + 25)
---
m + 25 = (3/5)(f + 25)
m + 25 = (3/5)f + 15
(1/3)f + (10/3) + 25 = (3/5)f + 15
(10/3) + 10 = (3/5)f - (1/3)f
(10/3) + (30/3) = (9/15)f - (5/15)f
(40/3) = (4/15)f
---
f = (40/3)(15/4)
f = (10)(5)
f = 50
---
m = (1/3)(f - 5) + 5
m = (1/3)(50 - 5) + 5
m = (1/3)(45) + 5
m = 15 + 5
m = 20
---
answer:
m = 20
f = 50
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php